OpenStudy (anonymous):

Find the midpoint between two points on a coordinate plane whose coordinates are (5, 7) and (-3, 1). (2, 8) (1, 4) (8, 8) (4, 4)

OpenStudy (hoblos):

the coordinates of the midpoint are x = (x1+x2)/2 y= (y1+y2)/2 so here x= (5-3)/2 = 1 y= (7+1)/2 = 4 thus the midpoint is (1,4)